# A relative version of Bass’ theorem about finite-dimensional algebras

### v1

@@ 3. Flaprojective Modules
Let $R$ be an associative ring. It is easy to prove that the pair of classes of flat left $R$-modules and cotorsion left $R$-modules $(\mathsf{F},\mathsf{C})$ $=$ $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ is a cotorsion pair in $R{\operatorname{\mathsf{--Mod}}}$[16, Lemma 3.4.1], [13, Lemma 1.3]. Proving that this pair of classes, called the flat cotorsion pair or the Enochs cotorsion pair [8], [9, Definition 5.18(i)]), is a complete cotorsion pair, is harder. There are two proofs of this result in the paper [3], one of them based on the following observation of Enochs.
@@
This is [3, Lemma 1 and Proposition 2]. See [9, Lemma 6.17] for a generalization. ∎
@@
For any associative ring $R$, the flat cotorsion pair $(\mathsf{F},\mathsf{C})$ $=$ $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ in $R{\operatorname{\mathsf{--Mod}}}$ is complete.
@@
This is [3, Section 2] and a particular case of [9, Theorem 6.19]. The assertion follows from Proposition 3.1 by virtue of Corollary 1.3(b). ∎
Now let $R\longrightarrow A$ be a homomorphism of associative rings. A left $A$-module $F$ is said to be flaprojective relative to $R$ (or $A/R$-flaprojective) [13, Section 1.6] if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion as left $R$-modules (i. e., the underlying left $R$-module of $C$ is cotorsion).
Since the class of cotorsion $R$-modules is closed under cokernels of monomorphisms (see the beginning of Section 2), so is the class of all $A$-modules that are cotorsion over $R$. All cotorsion $A$-modules are cotorsion over $R$ by Lemma 2.3; in particular, all injective $A$-modules are cotorsion over $R$.
Using these observations, one can show that $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all $A/R$-flaprojective left $R$-modules $F$, all left $A$-modules $C$ that are cotorsion over $R$, and all integers $n\geq 1$ (cf. [9, Lemma 5.24] and [14, Lemma 7.1]). It follows that the class of $A/R$-flaprojective $A$-modules is closed under kernels of epimorphisms in $A{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite direct sums. By Lemma 1.1, the class of $A/R$-flaprojective left $A$-modules is closed under transfinitely iterated extensions.
@@
(1) Since all cotorsion $A$-modules are cotorsion over $R$, it follows that all $A/R$-flaprojective $A$-modules are flat [13, Remark 1.28]. Obviously, all projective $A$-modules are $A/R$-flaprojective; so the class of $A/R$-flaprojective $A$-modules is, generally speaking, intermediate between the classes of projective and flat $A$-modules.
(2) If $A=R$ and $R\longrightarrow A$ is the identity ring homomoprhism, then the class of $A/R$-flaprojective $A$-modules coincides with the class of flat $R$-modules [13, Example 1.27(1)].
@@
This is a particular case of [5, Proposition VI.4.1.3]. See also [10, formula (1.1) in Section 1.2] or [13, Lemma 1.1]. ∎
@@
(a) the pair of classes of left $A$-modules $(\mathsf{F},\mathsf{C})$ $=$ $(A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}})$ is a complete cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$;
@@
This is [12, Proposition 3.3]. Notice that the left $A$-module $A^{+}=\operatorname{Hom}_{\mathbb{Z}}(A,\mathbb{Q}/\mathbb{Z})$ is injective, hence pure-injective, hence cotorsion; so by Lemma 2.3,  $A^{+}$ is a cotorsion left $R$-module and [12, Proposition 3.3] is applicable.
@@
We call the complete cotorsion pair $(\mathsf{F},\mathsf{C})$ $=$ $(A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}})$ from Theorem 3.5 the flaprojective cotorsion pair or the $A/R$-flaprojective cotorsion pair in the category of left $A$-modules $A{\operatorname{\mathsf{--Mod}}}$.
@@
(3) $\Longleftrightarrow$ (4) holds because all $A/R$-flaprojective $A$-modules are always flat by Remark 3.3(1).
@@
However, this assertion can be proved in a greater generality with the classical methods. The projectivity condition on the right $R$-module $A$ can be dropped here; it suffices to assume that $A$ is a finitely generated right $R$-module. Indeed, by [1, Theorem P(6)], a ring $S$ is left perfect if and only if all descending chains of cyclic right $S$-modules terminate. According to [4, Theorem 2], this holds if and only if all descending chains of finitely generated right $S$-modules terminate. It remains to observe that, whenever $A$ is a finitely generated right $R$-module, a right $A$-module is finitely generated if and only if it is finitely generated as an $R$-module.
@@
Of course, neither Theorem 2.8 nor Corollary 3.7 are true without the condition that the right $R$-module $A$ is finitely generated. If this condition is dropped, it suffices to consider the case when $R=k$ is a field and $A=k[t]$ is the $k$-algebra of polynomials in one variable $t$. Then the $A$-module $k(t)$ of rational functions in $t$ and the $A$-module $k[t,t^{-1}]$ of Laurent polynomials in $t$ are flat but not projective. So there exist $A$-modules that are not cotorsion (in fact, the free $A$-module $A$ is not cotorsion). However, every $k$-vector space is a cotorsion $k$-module, so all $A$-modules are cotorsion over $k$. Therefore, all $A/k$-flaprojective $A$-modules are projective.
However, we do not know whether the assertions of Theorem 2.8 and/or Corollary 3.7 hold true with the projectivity assumption on the right $R$-module $A$ dropped. For example, if $A$ is an associative algebra over a commutative ring $R$ and the $R$-module $A$ is finitely presented, does it follow that the ring homomorphism $R\longrightarrow A$ is relatively (left) perfect?

### v2

@@ 3. Flaprojective Modules
Let $R$ be an associative ring. It is easy to prove that the pair of classes of flat left $R$-modules and cotorsion left $R$-modules $(\mathsf{F},\mathsf{C})$ == $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ is a cotorsion pair in $R{\operatorname{\mathsf{--Mod}}}$[18, Lemma 3.4.1], [14, Lemma 1.3]. Proving that this pair of classes, called the flat cotorsion pair or the Enochs cotorsion pair [8], [10, Definition 5.18(i)]), is a complete cotorsion pair, is harder. There are two proofs of this result in the paper [3], one of them based on the following observation of Enochs.
@@
This is [3, Lemma 1 and Proposition 2]. See [10, Lemma 6.17] for a generalization. ∎
@@
For any associative ring $R$, the flat cotorsion pair $(\mathsf{F},\mathsf{C})$ == $(R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$, $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}})$ in $R{\operatorname{\mathsf{--Mod}}}$ is complete.
@@
This is [3, Section 2] and a particular case of [10, Theorem 6.19]. The assertion follows from Proposition 3.1 by virtue of Corollary 1.3(b). ∎
Now let $R\longrightarrow A$ be a homomorphism of associative rings. A left $A$-module $F$ is said to be flaprojective relative to $R$ (or $A/R$-flaprojective) [14, Section 1.6] if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion as left $R$-modules (i. e., the underlying left $R$-module of $C$ is cotorsion).
Since the class of cotorsion $R$-modules is closed under cokernels of monomorphisms (see the beginning of Section 2), so is the class of all $A$-modules that are cotorsion over $R$. All cotorsion $A$-modules are cotorsion over $R$ by Lemma 2.3; in particular, all injective $A$-modules are cotorsion over $R$. Using these observations, one can show that $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all $A/R$-flaprojective left $R$-modules $F$, all left $A$-modules $C$ that are cotorsion over $R$, and all integers $n\geq 1$ (cf. [10, Lemma 5.24] and [15, Lemma 7.1]).
It follows that the class of $A/R$-flaprojective AA-modules is closed under kernels of epimorphisms in $A{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite direct sums. By Lemma 1.1, the class of $A/R$-flaprojective left $A$-modules is closed under transfinitely iterated extensions.
@@
(1) Since all cotorsion $A$-modules are cotorsion over $R$, it follows that all $A/R$-flaprojective AA-modules are flat [14, Remark 1.29]. Obviously, all projective $A$-modules are $A/R$-flaprojective; so the class of $A/R$-flaprojective AA-modules is, generally speaking, intermediate between the classes of projective and flat $A$-modules.
(2) If $A=R$ and $R\longrightarrow A$ is the identity ring homomoprhism, then the class of $A/R$-flaprojective AA-modules coincides with the class of flat $R$-modules [14, Example 1.28(1)].
@@
This is a particular case of [5, Proposition VI.4.1.3]. See also [11, formula (1.1) in Section 1.2] or [14, Lemma 1.1]. ∎
@@
(a) the pair of classes of left $A$-modules $(\mathsf{F},\mathsf{C})$ == $(A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}})$ is a complete cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$;
@@
This is [13, Proposition 3.3]. Notice that the left $A$-module $A^{+}=\operatorname{Hom}_{\mathbb{Z}}(A,\mathbb{Q}/\mathbb{Z})$ is injective, hence pure-injective, hence cotorsion; so by Lemma 2.3,  $A^{+}$ is a cotorsion left $R$-module and [13, Proposition 3.3] is applicable.
@@
We call the complete cotorsion pair $(\mathsf{F},\mathsf{C})$ == $(A{\operatorname{\mathsf{--Mod}}}_{A/R{\operatorname{\mathsf{-flpr}}}}$, $A{\operatorname{\mathsf{--Mod}}}^{R{\operatorname{\mathsf{-cot}}}})$ from Theorem 3.5 the flaprojective cotorsion pair or the $A/R$-flaprojective cotorsion pair in the category of left $A$-modules $A{\operatorname{\mathsf{--Mod}}}$.
@@
(3) $\Longleftrightarrow$ (4) holds because all $A/R$-flaprojective AA-modules are always flat by Remark 3.3(1).
@@
However, this assertion can be proved in a greater generality with the classical methods. The projectivity condition on the right $R$-module AA can be dropped here; it suffices to assume that $A$ is a finitely generated right $R$-module. Indeed, by [1, Theorem P(6)], a ring $S$ is left perfect if and only if all descending chains of cyclic right $S$-modules terminate. According to [4, Theorem 2], this holds if and only if all descending chains of finitely generated right $S$-modules terminate. It remains to observe that, whenever $A$ is a finitely generated right $R$-module, a right $A$-module is finitely generated if and only if it is finitely generated as an $R$-module.
@@
Of course, neither Theorem 2.8 nor Corollary 3.7 are true without the condition that the right $R$-module AA is finitely generated. If this condition is dropped, it suffices to consider the case when $R=k$ is a field and $A=k[t]$ is the $k$-algebra of polynomials in one variable $t$. Then the $A$-module $k(t)$ of rational functions in $t$ and the $A$-module $k[t,t^{-1}]$ of Laurent polynomials in $t$ are flat but not projective. So there exist $A$-modules that are not cotorsion (in fact, the free $A$-module AA is not cotorsion). However, every $k$-vector space is a cotorsion $k$-module, so all $A$-modules are cotorsion over $k$. Therefore, all $A/k$-flaprojective AA-modules are projective.
However, we do not know whether the assertions of Theorem 2.8 and/or Corollary 3.7 hold true with the projectivity assumption on the right $R$-module AA dropped. For example, if $A$ is an associative algebra over a commutative ring $R$ and the $R$-module AA is finitely presented, does it follow that the ring homomorphism $R\longrightarrow A$ is relatively (left) perfect?
It is clear from Corollary 3.6(1) or (2) that the composition of any two relatively left perfect homomorphisms of associative rings is a relatively left perfect homomorphism of associative rings again. The following example describes another sufficient condition for a homomorphism of associative rings to be relatively left perfect.
Example 3.10.
This example is due to S. Bazzoni and is reproduced here with her kind permission.
A homomorphism of associative rings $R\longrightarrow A$ is said to be a ring epimorphism if it is an epimorphism in the category of associative rings, or equivalently, if the functor of restriction of scalars $A{\operatorname{\mathsf{--Mod}}}\longrightarrow R{\operatorname{\mathsf{--Mod}}}$ is fully faithful. A ring homomorphism $R\longrightarrow A$ is an epimorphism if and only if the natural map of abelian groups $N\otimes_{R}M\longrightarrow N\otimes_{A}M$ is an isomorphism for every right $A$-module NN and left $A$-module $M$. A ring epimorphism $R\longrightarrow A$ is said to be left flat if $A$ is a flat left $R$-module. We suggest the book [17, Sections XI.1–3] and the paper [9, Section 4] as the reference sources on (left flat and arbitrary) ring epimorphisms.
Let $R\longrightarrow A$ be a left flat ring epimorphism and $F$ be a left $A$-module. Then the following conditions are equivalent:
  1. (1)
    $F$ is flat as an $R$-module;
  2. (2)
    $F$ is flat as an $A$-module;
  3. (3)
    $F$ is an $A/R$-flaprojective AA-module.
Indeed, every flat left $A$-module is flat as an $R$-module since $A$ is a flat left $R$-module. Conversely, for any left $A$-module $F$ we have isomorphisms of left $A$-modules $F\simeq A\otimes_{A}F\simeq A\otimes_{R}F$, since $R\longrightarrow A$ is a ring epimorphism. Now if $F$ is flat as an $R$-module, then it is clear that $F\simeq A\otimes_{R}F$ is flat as an $A$-module.
Furthermore, for any homomorphism of associative rings, every $A/R$-flaprojective AA-module is flat as an $A$-module by Remark 3.3(1). Conversely, by Corollary 3.6, in order to show that all flat left $A$-modules are $A/R$-flaprojective, it suffices to check that all left $A$-modules that are cotorsion over $R$ are also cotorsion over $A$.
Let $C$ be an $R$-cotorsion left $A$-module. For any left $A$-module $C$ we have isomorphisms of left $A$-modules $C\simeq\operatorname{Hom}_{A}(A,C)\simeq\operatorname{Hom}_{R}(A,C)$, since $R\longrightarrow A$ is a ring epimorphism. Now the left $R$-module AA is flat by assumption, while $C$ is a cotorsion left $R$-module. Applying Lemma 2.4 for the ring $S=A$ and the $R$-AA-bimodule $K=A$, we conclude that the left $A$-module $C\simeq\operatorname{Hom}_{R}(A,C)$ is cotorsion.
Thus all left flat epimorphisms of associative rings are relatively left perfect as associative ring homomorphisms.